Hawaii-2019-03 experiment with MOBYrefresh - BS04cfg101 & RS03cfg002 - laser data


Day [1 2]
BS04cfg101 = barefiber + back from Andor/Resonon after window replacement
RS03cfg002 = barefiber


Problems seen (and solved) and good things:

  1. UV ghost is still being seen. See Page Num 4.01. Interestingly only 2 of the LEDs had the UV ghost which I think improves the possibility that it can be characterized.

Log Sheets BS04/RS03

Photos

Track pixels:   BS04 day 1 data
Track:   1   2   3   4   5   6   7   8   9  10  11  12  13   14 
Left:    1  75 150 223 296 369 441 512 585 657 729 800 874  949 
Right:  74 149 222 295 368 440 511 584 656 728 799 873 948 1024 
Track pixels:   RS03 Day 1
Track:   1   2   3   4   5   6   7   8   9  10  11  12  13   14  
Left:    1  81 156 229 301 373 445 516 588 659 732 804 878  953  
Right:  80 155 228 300 372 444 515 587 658 731 804 877 952 1024  
Laser wavelength calibration:

P(1,:) = [-1.035002e-08 2.725833e-05 0.3312411 345.0497];
P(2,:) = [-1.13027e-08 2.846401e-05 0.3308512 345.023];
P(3,:) = [-1.16374e-08 2.887411e-05 0.3307137 344.9965];
P(4,:) = [-1.218432e-08 2.955162e-05 0.3304919 344.9896];
P(5,:) = [-1.286719e-08 3.043082e-05 0.3301905 344.9923];
P(6,:) = [-1.362072e-08 3.142773e-05 0.329827 345.0249];
P(7,:) = [-1.404966e-08 3.198142e-05 0.32962 345.0305];
P(8,:) = [-1.454766e-08 3.262885e-05 0.3293898 345.0223];
P(9,:) = [-1.462032e-08 3.279386e-05 0.3292854 344.9869];
P(10,:) = [-1.462348e-08 3.286702e-05 0.3292309 344.9555];
P(11,:) = [-1.451556e-08 3.277941e-05 0.3292396 344.9159];
P(12,:) = [-1.492203e-08 3.32566e-05 0.3291853 344.928];
P(13,:) = [-1.504552e-08 3.340387e-05 0.3291913 344.9195];
P(14,:) = [-1.449819e-08 3.269468e-05 0.3294432 344.8761];

Laser wavelength calibration:

P(1,:) = [-1.915839e-08 4.711645e-05 0.3698228 503.0073];
P(2,:) = [-1.659526e-08 4.195416e-05 0.3729314 502.4745];
P(3,:) = [-1.511054e-08 3.89291e-05 0.3747683 502.1402];
P(4,:) = [-1.422174e-08 3.704621e-05 0.3759458 501.9213];
P(5,:) = [-1.408619e-08 3.670112e-05 0.3761747 501.8664];
P(6,:) = [-1.359309e-08 3.564731e-05 0.3768463 501.7464];
P(7,:) = [-1.400264e-08 3.64829e-05 0.3763156 501.8302];
P(8,:) = [-1.312091e-08 3.476771e-05 0.3773297 501.6189];
P(9,:) = [-1.229404e-08 3.319167e-05 0.378253 501.4104];
P(10,:) = [-1.007216e-08 2.882861e-05 0.380908 500.8817];
P(11,:) = [-7.289769e-09 2.332863e-05 0.384284 500.2337];
P(12,:) = [-6.459531e-09 2.176121e-05 0.38526 499.9816];
P(13,:) = [-1.79322e-08 4.488718e-05 0.3710214 502.5023];
P(14,:) = [-1.789332e-08 4.477433e-05 0.3712088 502.3902];

All the raw data files and their images , Table of KEYWORDS variable
Page Number
Link
Description
Date
1.01
Day 01 - 17 Jul Raw Data: BS & RS | Table : BS & RS
DAY01 - all track LED check
17 Jul 2019
1.02
Day 02 - 18 Jul Raw Data: BS & RS | Table : BS & RS
DAY02 - all track laser data - bare fiber
18 Jul 2019
1.03
Day 03 - 19 Jul Raw Data: BS & RS | Table : BS & RS
DAY03 - Track 6 Colored Source ThorLabs LEDs
19 Jul 2019
1.04
Day 04 - 26 Aug Raw Data: BS & RS | Table : BS & RS
DAY04 - Track 6 constant wavelength (cw) laser diodes @ 407 nm sat and unsat
27 Aug 2019
 
Day 1
2.01 track definitions BS/ RS Getting the new track definitions  
Day 2
3.01 All tracks laser data  
3.02 laser wavecal BS T1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14 BSG Laser Wavelength cal  
3.03 laser wavecal RS T 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14 RSG Laser Wavelength cal  
3.04 wavecal comp BS/ RS Compare the wavelength calibration before and after the window repair  
3.05 Ambient Check BS / RS Light check using Ambient data set  
Day 3
4.01 UV ghost BS/ RS UV ghost showing up in some of the LED data sets  
Day 4
5.01    
Compiled data sets
30.01 Daily Track Movement -> BS link & RS link Laser Track movelength graphs for all days data 22 Jan 2019
30.03 backgrounds-> BS link & RS link Backgrounds from all experments  
30.04 Low Pixel nets-> BS link & RS link Mike asked why the RS Low pixel (1-90) data are negative (so I made these graphs)  
30.05 All laser data-> BS link & RS link All the laser data by track collected so far  
DAY 1 EMAIL from Mike

Hi Steph,
Riley & I managed to collect some BS04cfg101 & RS03cfg001 data yesterday,
which are at /ftp1/Mike/HI-2019-03/ = ready for your data processing help.

I wanted to setup both specs to collect data somewhat-simultaneously -
I'm not sure that doing one-spec-at-a-time was such a bad idea now...

There are photos of the setup, in the cal van next to the tent.
Both specs' bare fiber-bundle fibers were looking into the Spectralon sphere.
There were no internal aux data (degC, %RH) collected from the specs.

Day01 had as input to the sphere a Thorlabs LED.  I think this was the first use
of this source.  Carol found this in 2015 and we bought one = LED driver /
power supply, aka Thorlabs DC2100, plus 5x different LED "heads". Last night
we scanned the M530L3 head, which should be centered ~ 530 nm.
The LED pwr sup was set at 9 mA ( == LOW power ), as noted in the log sheet.

I ran the Labsphere Si PhotoDiode monitor at the sphere monitor port #A.
This would be the PD that Carol had at NIST for cals (which you should have
data from?) but, the detector was not a UV-enhanced kine, so it has no
response in the UV, and we will try to re-purchase another PD with a
UV-enhanced detector.  The PD's electrical-current to voltage pre-amp was
the one Carol had at NIST for the NIST cals, aka Femto DLPCA-200. The Voltage
output from the pre-amp was logged via the 34970A DVM using software
Vdc6_5.vi, which you have seen output from before.  Someday I will update
this VI for use with a PD monitor...  Last night, the PD preamp went to DVM
chan#3 (i.e. setup for high precision Voltage measurement, typically for a
shunt during Ed/Lu absolute lamp cals),  and DVM chan#5 (mike said this is really chan#6) had the Radio Shack
10 kOhm thermistor for ambient temperature monitor. For now the only
indication of what gain setting the PD preamp was set at is noted in the
log sheet - last night was @ low gain, 10^4.  This will someday be included
in the log file's header...

The log file from the DVM is in the HI-2019-03/BSRS/ subdir on the NAS.
The 1x day01 log sheet is in the HI-2019-03/doc/ subdir on the NAS.

One problem last night: I did not power ON the TEC for the PD monitor.
Tonight I will try harder to turn that bugger on.

We had to re-load the Andor SOLIS software onto 2x HP laptops to control the
BS & RS.  So I'm hoping that I set all acquisition parameters correctly.  The data
acq sequence was 2x bac, 4x sig, 2x bac.

Tonight we will attempt to collect some Ekspla laser scans via BS & RS with
all bare bundle fibers illuminated.

Good Luck with this!  MF

Part 2 email

Hi Steph,
Rodger that.  Thanks for catching this!
Was the cfg002 bare-FO bundle kine in Miami?
I know of no cfg changes since the RS03's return to HI,
so maybe we should stay stuck on cfg002?
The RS03 did not get any "major" update (ex. wedged window)
so I don't think we should go to cfg101.

I had just written here a sentence to ax you for a cfg page, then I found I way
to "MOBYrefresh - Configurations", which answered I question,  butt,
it looks like there is no listing for cfg002 for RS03 there...

However, I do see under "MOBYrefresh - RS03 Configurations" page
that there was a cfg002 for HI-2016-05, but the one-and-only log sheet says cfg001.
So I think it should be RS03cfg001 for HI-2016-05 (?).

And, since there was no log sheet for Miami-2016-02 (oh my!) there must have
been some reason to go from cfg001 to cfg002 (?).  Maybe the cfg change was
tightening the screws that came loose during shipping to Miami?

In conclusion:
I just updated the HI-2019-03 day01 log sheet to say RS03cfg002,
and, I FTPed it to NAS /ftp1/Mike/HI-2019-03/doc/,
and, I changed the name of this file because it was of non-standard format.
MF

DAY 2 EMAIL from Mike
Hi Steph,
HI-2019-03 day02 data are heading to the NAS.
Day02 was Ekspla laser scans via all 28x bundle fibers.

Mostly the same instrument setup as day01,
except I re-did the black-cloth over the FO bundles and
over the sphere exit port & jig, after taking photos last night.
And I remembered to power on the PD monitor's TEC tonight!

At the very start, with no laser input to the integrating sphere,
I did an ambient scan set to check if the van lights were
getting into the BS & RS - these would be bac & sig scans #001:004.
The van lights were left on during all scans on day01 & day02.

I choose some curious wavelengths - 455,530,617,625,730 nm -
because they are centre waves for the Thorlabs LEDs we have.

Then I chose some waves to fill out to the ends of the BS & RS.
The 350 nm line showed the second order signal on the BS.
And 345 nm was just-off the short end of the BS, but the 2nd order was on the BS.
I am very interested to see if the possible "sphere fluorescence" shows up via
the 350 and 345 nm lines.

I had to keep changing the laser power to get on-scale scans,
so the pump laser UV mJ readings are noted on the cal log.
It will be interesting to see how the mJ relate to the PD monitor Voltages...
I did not try changing the PD's preamp gain during these scans.

My hope is to run Ekspla again tomorrow with only 1x FO from each spec,
then Friday run 1x FO-each scans of the 5x Thorlabs LEDs.  I suppose this
plan might change after the telecon tomorrow!
MF

DAY 3 EMAIL from Mike
Hi Steph,
The Expla laser decided to take the day off.
We spent some time trying to coax it back to work, butt to no avail.
Hopefully the Ekspla will decide to work tomorrow...

So today we ran the 5x Thorlabs LED sources via only Trk#6 = ON
on the BS04 and RS03.  There are 5x new PD mon files for these LEDs
( plus 1x PD mon file from the Ekspla when it was having problems ).
MF

DAY 3 continued EMAIL from Mike
Thanks Steph, I think...
I think your pg. 4.01 = "UV ghost BS", does show an interesting situation!

I tried to well-cover with black-tape all the non illuminated fiber tips, so I don't have much
hope that there was a simple light leak into not-Trk6-fibers from the now-open holes in the jig
at the sphere exit port.

I am sorry to say that the Ekspla laser seems to have died - yesterday and again today -
so there were no single track laser data from today / day04.
I'll have to contact tech support once I return from vacation.

Good luck scratching your heads over this one...
MF
DAY 4 continued EMAIL from Mike
Hello Steph,
I finally got back to making more BS04+RS03 scans under the HI-2019-03 kine.
There are some day04 data for you at the NAS /ftp1/Mike/HI-2019-03/
= one data sheet + 12x pairs of bac & sig .fits files from the BS & RS + three photos.

All scans sets are 2 bac, 4 sig, 2 bac.  Only trk#6 illuminated.

The Ekspla tunable laser is not working, so I dug out the ole Power Technologies
constant wavelength (cw) laser diodes @ 372 & 407 nm.  The 407nm is too bright -
maybe I'll try dimming it via a ND filter (?).  So the 407 scans were saturating
the BS @ 35 msec exposure time (i.e. same 35ms as the shutter open/close time).

I also did a set of ambient light leak scans with the van lights = on.
This was to check the coupling of the laser to the sphere, and to check
that the not-illuminated fibers were properly dark ( iUsed a new technique
for darkening the unused fibers, suggested by Riley, shown in a photo).
However, iManaged to not collect the 2x bac scans after the 4x sig scans,
so iCopied the 2x bac scans from before the 4x sig scans to fill in this gap.

There are 2x PD mon files at /ftp1/Mike/HI-2019-03/BSRS/
one each for the 372 & 407 nm laser diodes.
I think these should be OK log files.
The PD setup with the Femto preamp should be OK with cw lasers.

I've been trying to understand the PD mon files from the Ekspla scans.
It turns out they are bogus.  One cannot monitor a pulsed laser via
a PhotoDiode + a current amplifier (aka preamp).  One must use one's
Electrometer and measure accumulated charge.

Aloha, MF